# Insulating and metallic phases in the one-dimensional Hubbard-Su-Schrieffer-Heeger model: Insights from a backflow-inspired variational wave function

### v2

@@ III.3 Results away from half-filling
We now move to the most important part of this work and examine the behavior of the ground state in the doped regime, with a number of electrons $N_{\rm e}<L$, and a doping defined by $\delta=1-N_{\rm e}/L$. We restrict our attention to the subspace where the total spin projection $\sum_{i}\hat{S}^{z}_{i}=0$. Let us start by doping the Mott insulator. In the Hubbard model, with only nearest-neighbor hopping $t$ and without phonons, a Luttinger liquid is obtained, with gapless modes in both charge and spin sectors Giamarchi (2003). In turn, a power-law decay (in real space) and cusps at $q=2k_{F}$ (in momentum space), where $k_{F}=\pi(1-\delta)/2$, are present in the ground-state correlations. These features are also observed in the presence of phonons, as long as the Mott insulator is stabilized at half-filling, see Fig. 4. Here, the cusps at $q=2k_{F}$ are clearly visible in the variational calculations of $N(q)$ and $S(q)$. We notice that cusps are less evident in the density correlations within DMRG, especially when approaching half-filling. Although these singularities are less pronounced than in the backflow wave function, the linear behavior of $N(q)$ for small values of $q$ (indicating gapless charge modes for $q\to 0$) leaves no doubt on the metallic nature of the system. We notice that, in the doped case, DMRG calculations are much more demanding than in the half-filled regime; nevertheless, the variance of the total energy is always lower than $0.09t^{2}$.
@@
The most interesting outcome appears when doping the Peierls insulator. In this case, a Luther-Emery liquid emerges at small values of the doping. Indeed, for $\delta>0$ the system immediately turns into a metal, but the spin gap remains finite close to half-filling. By further increasing $\delta$, the spin gap gradually decreases, eventually leading to a Luttinger liquid for sufficiently large values of $\delta$. This fact is particularly interesting since the Luther-Emery state has been found in repulsive models with no phonons, emerging from a spin-gapped and dimerized insulator Fabrizio (1996); Daul and Noack (1998); Nishimoto et al. (2008); Balents and Fisher (1996); Shen et al. (2023), however, its stabilization has been argued to heavily rely on the existence of multiple Fermi points in the non-interacting band structure. By contrast, in the present case the band structure is trivial, with only two Fermi points, and the spin gap is opened by the e-ph coupling. The results of the density and spin structure factors are reported in Fig. 5, where a smooth $S(q)$ persists in the doped system, before the insurgence of cusps at $q=2k_{F}$. The existence of a spin-gapped metal is confirmed by DMRG. Curiously, the Luther-Emery in the Hubbard-SSH model has not been detected before, even though its stability extends down to the $U=0$ limit, where a finite e-ph interaction drives into a Peierls state at half-filling, see Fig. 2. The exemplification of this fact is shown for $\lambda=0.16$ and $\delta=0.08$ in Fig. 6. Even though a precise determination of the power-law decay of correlation functions is difficult (given their oscillatory behavior), we have evidence that the bond-bond correlations are the dominant ones in the Luther-Emery phase (not shown).
@@ IV Conclusions
In conclusion, we showed that a variational Ansatz, which goes beyond the simple Jastrow-Slater approximation, can characterize the various phases of the one-dimensional Hubbard-Su-Schrieffer-Heeger model both at half-filling and in the hole-doped regime. Our results are in good agreement with DMRG calculations, confirming that the physical description of the system we are providing is accurate. In addition, we are able to convincingly show the presence of a spin-gapped metallic phase upon doping the Peierls insulator. Since the variational wave function we propose can be readily generalized in two dimensions and variational Monte Carlo methods do not suffer limitations from considering higher-dimensional models, our next step will be to study a two-dimensional Hubbard-Su-Schrieffer-Heeger model within the same framework. We are particularly interested in the doped regime, to investigate how the presence of phonons can favor various types of instabilities, such as stripe order or $d$-wave superconductivityWang et al. (2022). We remark that a one-dimensional band structure possesses nesting at $2k_{F}$, which makes the superconducting fluctuations compete against the bond-order ones, with the latter dominant in the Hubbard-Su-Schrieffer-Heeger model. In higher dimensions, nesting is a rather exceptional circumstance, especially away from commensurate densities, suggesting that, should our findings extend beyond one dimension, superconductivity would be the dominant instability channel.

### v3

@@ III.3 Results away from half-filling
We now move to the most important part of this work and examine the behavior of the ground state in the doped regime, with a number of electrons $N_{\rm e}<L$, and a doping defined by $\delta=1-N_{\rm e}/L$. We restrict our attention to the subspace where the total spin projection $\sum_{i}\hat{S}^{z}_{i}=0$. Let us start by doping the Mott insulator. In the Hubbard model, with only nearest-neighbor hopping $t$ and without phonons, a Luttinger liquid is obtained, with gapless modes in both charge and spin sectors [21]. In turn, a power-law decay (in real space) and cusps at $q=2k_{F}$ (in momentum space), where $k_{F}=\pi(1-\delta)/2$, are present in the ground-state correlations. These features are also observed in the presence of phonons, as long as the Mott insulator is stabilized at half-filling, see Fig. 4. Here, the cusps at $q=2k_{F}$ are clearly visible in the variational calculations of $N(q)$ and $S(q)$. We notice that cusps are less evident in the density correlations within DMRG, especially when approaching half-filling. Although these singularities are less pronounced than in the backflow wave function, the linear behavior of $N(q)$ for small values of $q$ (indicating gapless charge modes for $q\to 0$) leaves no doubt on the metallic nature of the system. We notice that, in the doped case, DMRG calculations are much more demanding than in the half-filled regime; nevertheless, the variance of the total energy is always lower than $0.09t^{2}$.
@@
The most interesting outcome appears when doping the Peierls insulator. In this case, a Luther-Emery liquid emerges at small values of the doping. Indeed, for $\delta>0$ the system immediately turns into a metal, but the spin gap remains finite close to half-filling. By further increasing $\delta$, the spin gap gradually decreases, eventually leading to a Luttinger liquid for sufficiently large values of $\delta$. This fact is particularly interesting since the Luther-Emery state has been found in repulsive models with no phonons, emerging from a spin-gapped and dimerized insulator [23, 24, 25, 26, 27], however, its stabilization has been argued to heavily rely on the existence of multiple Fermi points in the non-interacting band structure. By contrast, in the present case the band structure is trivial, with only two Fermi points, and the spin gap is opened by the e-ph coupling. The results of the density and spin structure factors are reported in Fig. 5, where a smooth $S(q)$ persists in the doped system, before the insurgence of cusps at $q=2k_{F}$. The existence of a spin-gapped metal is confirmed by DMRG. Curiously, the Luther-Emery in the Hubbard-SSH model has not been detected before, even though its stability extends down to the $U=0$ limit, where a finite e-ph interaction drives into a Peierls state at half-filling, see Fig. 2. The exemplification of this fact is shown for $\lambda=0.16$ and $\delta=0.08$ in Fig. 6. Even though a precise determination of the power-law decay of correlation functions is difficult (given their oscillatory behavior), we have evidence that the bond-bond correlations are the dominant ones in the Luther-Emery phase (not shown).
Figure 7: Phase diagram at $U/t=4$ as a function of $\lambda$ and $\delta$ with $\hbar\omega/t=1$. Calculations are performed on a chain with $L=50$ sites, with data points represented by filled circles. The system is insulating at $\delta=0$, indicated by the shaded region at the bottom of the figure. For finite values of $\delta$, a transition is observed from a Luttinger liquid to a Luther-Emery metallic phase.
Figure 8: Density, spin, and bond structure factors for a Luther-Emery metal at $\lambda=0.25$ and $\delta=0.15$; the quantities are shown for chains of lengths $L=40$, $80$, and $120$. No significant changes are observed in the spin structure factors, confirming that the Luther-Emery liquid survives in the large-chain limit; the peaks of the charge and bond structure factors at $2k_{F}$ diverge sub-linearly with the system size.
By analyzing the presence of singularities in the spin structure factor, we are able to construct the phase diagram where the electron doping $\delta$ is varied. In particular, we report the case in which $\lambda$ is varied and $U/t=4$ is kept fixed in Fig. 7. Here, a Luther-Emery metal exists when (slightly) doping the Peierls insulator at $\delta=0$ and sufficiently large values of $\lambda$. By further increasing the doping concentration, the ground state becomes a Luttinger liquid, with gapless spin excitations. To confirm the stability of the Luther-Emery phase and ensure that the spin gap persists as the system size increases, we also perform a size-scaling analysis, focusing at $\lambda=0.25$ and $\delta=0.15$ (which falls within the Luther-Emery phase phase) and display the structure factors for three different sizes (i.e., $L=40$, $80$, and $120$), see Fig. 8. Remarkably, the spin structure factor does not show substantial modifications, keeping a completely smooth behavior, which implies the stability of the Luther-Emery phase also on large systems (and eventually in the thermodynamic limit). For $U/t=4$, we rule out the presence of phase separation in both the Luttinger and Luther-Emery phases within the range of $\lambda$ and $\delta$ shown in Fig. 7. This conclusion is clearly supported by DMRG simulations, which gives the correct (convex) behavior of the energy as a function of the electron doping; VMC results show a tiny region of phase separation close to half filling in the Luther-Emery phase, but this fact is primarily due to the different accuracy of VMC energies, which is very high in the Peierls insulator and slightly worse for small $\delta$ (still grasping the correct physical behavior).
@@ IV Conclusions
In conclusion, we showed that a variational Ansatz, which goes beyond the simple Jastrow-Slater approximation, can characterize the various phases of the one-dimensional Hubbard-Su-Schrieffer-Heeger model both at half-filling and in the hole-doped regime. Our results are in good agreement with DMRG calculations, confirming that the physical description of the system we are providing is accurate. In addition, we are able to convincingly show the presence of a spin-gapped metallic phase upon doping the Peierls insulator. Since the variational wave function we propose can be readily generalized in two dimensions and variational Monte Carlo methods do not suffer limitations from considering higher-dimensional models, our next step will be to study a two-dimensional Hubbard-Su-Schrieffer-Heeger model within the same framework. We are particularly interested in the doped regime, to investigate how the presence of phonons can favor various types of instabilities, such as stripe order or $d$-wave superconductivity[46]. We remark that a one-dimensional band structure possesses nesting at $2k_{F}$, which makes the superconducting fluctuations compete against the bond-order ones, with the latter dominant in the Hubbard-Su-Schrieffer-Heeger model. In higher dimensions, nesting is a rather exceptional circumstance, especially away from commensurate densities, suggesting that, should our findings extend beyond one dimension, superconductivity would be the dominant instability channel.